OpenStudy (anonymous):

The midpoint of Segment cd is E( -1,0). one end point is C(5,2). what are the other coordinates to the other end point??

OpenStudy (anonymous):

\[Midpoint = \left( \frac{ x_{1}+x _{2} }{ 2 }, \frac{ y_{1}+y _{2} }{ 2 } \right)\] So we know the midpoint is (-1,0). We can use (5,2) as (x1,y1).

OpenStudy (anonymous):

oh okay cause i was thinking i should use segment addition postulate equation